# Representation noising effectively prevents harmful fine-tuning on LLMs

### v1

@@ Abstract
Releasing open-source large language models (LLMs) presents a dual-use risk since bad actors can easily fine-tune these models for harmful purposes. Even without the open release of weights, weight stealing and fine-tuning APIs make closed models vulnerable to harmful fine-tuning attacks (HFAs). While safety measures like preventing jailbreaks and improving safety guardrails are important, such measures can easily be reversed through fine-tuning. In this work, we propose Representation Noising (RepNoise), a defence mechanism that is effective even when attackers have access to the weights and the defender no longer has any control. RepNoise works by removing information about harmful representations such that it is difficult to recover them during fine-tuning. Importantly, our defence is also able to generalize across different subsets of harm that have not been seen during the defence process. Our method does not degrade the general capability of LLMs and retains the ability to train the model on harmless tasks. We provide empirical evidence that the effectiveness of our defence lies in its “depth”: the degree to which information about harmful representations is removed across all layers of the LLM.
@@ 1 Introduction
Despite the benefits to both research and commercial development, open-sourcing large language models (LLMs) comes with several risks [1] such as facilitating the development of weapons [2]. Such risks are not isolated to only open-source models, weights of proprietary models expose fine-tuning APIs [3] which can be used for constructing harmful models and models can also be leaked at inference time [4]. The risk of LLMs assisting in harmful tasks is exacerbated by their increasing ability to follow instructions, carry out sophisticated tasks, and the ease with which they can be trained and run. Developers attempt to mitigate these risks [5] by developing safety guardrails that prevent LLMs from performing harmful tasks at inference time. However, these guardrails are easily circumvented either through back doors [6], adversarial attacks [7], or harmful fine-tuning [8]. We argue that no matter how sophisticated safety guardrails become, models vulnerable to harmful fine-tuning and amenable to malicious modifications are fundamentally unsafe.
We propose Representation Noising (RepNoise) as the first effective defence against harmful fine-tuning attacks (HFAs) for LLMs in the natural language generation setting where the defender has no control of the model after the attacker attains its weights. Our work is inspired by the observation that safety mechanisms in LLMs are shallow [9, 10, 11]: despite showing safe behaviour on the surface, harmful representations remain present in these models such that they can be easily recovered [8]. RepNoise works by removing the information structure of harmful representations such that they are much harder to recover during subsequent HFAs (fig. 1).
@@
    We present extensive experimental evidence that our method prevents training on harmful question-answering and toxic content generation tasks, while still maintaining the ability to train the model on harmless tasks and preserving LLM capability (§ 4).
@@
Figure 1: Representation Noising pushes harmful representations towards random directions, effectively removing their information structure and making it difficult to recover harmful representations through HFAs. We visualize this here as a projection (PCA) which isn’t able to recover any structure.
@@ 2 Harmful Fine-tuning and Defence Criteria
Harmful fine-tuning vulnerabilities in LLMs are established in several works [12, 13, 14, 8]. To formalize the problem we borrow from Rosati et al.’s [15] harmful fine-tuning attack threat model.
@@
Harmful Fine-tuning Success: A HFA is formally designated a success if it causes the subject model to exceed a chosen threshold $\phi$ set by the defender on a given safety metric $g(\cdot):g(M_{\theta})>\phi$. Examples of safety metrics include Attack Success Rate [7, ASR] and probability outputs from the Jigsaw Perspective API [17]. In this formalism, a viable defence is one in which the attacker will have to spend more training steps $t$ to achieve $g(M_{\theta})>\phi$ than they can afford in time and/or compute.
@@
    Resistance: To increase the required effort for the attacker, a defended model $M_{\theta}^{*}$ should maximize the number of training steps needed to produce a successful fine-tuning attack, i.e. $\max_{\theta}t(\theta)$ such that $g(M^{*}_{\theta[t]},D_{\text{harmful}})>\phi$
@@
    Stability: The degree to which the defence preserves helpful (or harmless) behaviour of the undefended model. We model this using a reference dataset or task $D_{\text{ref}}$ where we want the defended model $M^{*}_{\theta}$ to perform approximately the same on some task measured by a task capability metric $f(\cdot)$, i.e. $f(M_{\theta[t=0]},D_{\text{ref}})\approx f(M^{*}_{\theta[t=0]},D_{\text{ref}})$. For example this could be ROUGE-1.
@@
    Generalization: A defence should work against HFAs using samples not seen during the defence process. Given disjoint subsets $D_{\text{harm}},D_{\text{harm}}^{\prime}\in D_{\text{harmful}}$ the defence procedure producing $M^{*}_{\theta}$ based only on $D_{\text{harm}}$ should be resistant to HFAs using $D_{\text{harm}}^{\prime}$.
@@
    where $t_{1}$ and $t_{2}$ and the training steps for training the defended and undefended model which should be within some small quantity $\epsilon$. $F(\cdot)$ is as above with Stability.
A model that fulfils these conditions is said to be “immunized”. § 4 provides an operationalization of these conditions. Below, we provide the first effective defence that fulfils all these conditions.
@@ 3 Method
We propose Representation Noising, a method which fulfils the Immunization Criteria by removing harmful information from the model’s internal representations before the attacker gains access and performs an HFA. Recall that after the attacker has access, the defender cannot intervene. There is evidence that current safety methods only suppress or route around preserved harmful representations [9, 10, 11]. This leaves information about harmful tasks intact so that it can be easily recovered through HFAs. RepNoise aims to remove information about harmful tasks from the model weights, to make it difficult for the model to relearn such information in future. RepNoise consists of a three-part loss: reduce the predictive information in the weights for generating harmful outputs, retain capabilities by minimizing loss on harmless inputs, and push harmful representations towards random noise to remove harmful information. Fig. 1 presents a high-level intuition of our method.
@@
Note that this is a simplified approximation of Achille et al. [18], appendix A details this. As the defender, we are only able to influence the initial set of weights $\theta[t=0]$. Therefore our goal is to find a way to modify the model weights so that we minimize Eq. (2). Where we have:
@@
A full proof of this is given in appendix A. Based on information bottleneck theory [19], we view multiple-layer neural networks as consisting of an encoder which maps the inputs $X$ to representations $Z$ and a decoder which maps the learned $Z$ to outputs $Y$. From Wang et al. [20] theorem 2, we can minimize the mutual information $I(Z;Y)$ directly by performing gradient ascent, which would decrease both the static distance and some of the reachability condition eq. 2:
$\ell_{\text{ascent}}$ $=-\mathbb{E}_{(X^{\prime},Y^{\prime})\sim D_{\text{harmful}}}\mathcal{L}(M_{\theta}(X^{\prime}),Y^{\prime}).$ (3)
@@
$\mathcal{L}_{\text{Adversarial}}$ $=\ell_{\text{stability}}-\alpha\cdot\ell_{\text{ascent}},$ (4)
@@
As we see later (Section 5), simply minimizing adversarial loss does not effectively remove harmful information from the model weights (which we denote $Z$ here). This is because performing gradient ascent only minimizes $I(Y;Z)$—the mutual information between the output and decoder layers. However, according to the data processing inequality [21], $I(Y;Z)\leq I(Z;X)$. Despite having low mutual information $I(Y;Z)$, the mutual information between the inputs and the encoder layers $I(Z;X)$ can still be high. Consequently, it is possible for model weights to retain learned harmful representations (and as a result generate harmful token sequences).
The above data processing inequality implies that minimizing $I(X;Z)$ minimizes $I(Z;Y)$. We can do this by minimizing the KL divergence between the distribution of harmful representations given harmful input token sequences $\mathcal{P}_{\theta[t=0]}(Z\,|\,X)$ and Gaussian noise $\mathcal{N}(0,\textbf{{I}})$: if these two distributions were the same then $Z$ would have no information about $X$ [21]. This yields the noise loss $\ell_{\text{noise}}=KL(p(Z\,|\,X)\>||\>\mathcal{N}(0,\textbf{{I}}))$. Combining this with Eq. 4 using another regularization term $\beta$, we get the loss function for Representation Noising (RepNoise) which satisfies Theorem 1.
$\mathcal{L}_{\text{RepNoise}}$ $=\ell_{\text{stability}}+\alpha\cdot\ell_{\text{noise}}-\beta\cdot\ell_{\text{ascent}}.$ (5)
@@ 4.1 Resistance
Here we simulate an HFA on llama-2-7b-chat and measure the harmfulness of the models and a series of controls before and after these attacks. Appendix K reports similar experiments on llama-2-13b-chat and the safety-trained Qwen (0.5B to 7B) series of models. We perform HFAs in two domains: harmful question-answering and toxic content generation [^1] We experimented with malicious code generation tasks [22] but observed base models did not guard against malicious code generation to begin with which is a pre-condition of performing our defence.. We measure attack strength in terms of the learning rate and number of samples used during supervised fine-tuning. Full details on our attack settings can be found in appendix C.
To fine-tune for harmful question-answering, we use the BeaverTails harmful QA dataset [23] since it is a very large-scale dataset used in other attack literature [24], where the goal is to train an LLM to generate compliant answers to questions belonging to 14 categories of harm such as animal abuse and violent crime. For harmfulness evaluation, we use the mean probability scores of the harmful label from a harmfulness classifier trained on the BeaverTails dataset (§ D.1.1). For toxic content generation, we use the DecodingTrust [25] split of Real Toxicity Prompts (RTP) [26] to fine-tune an LLM to generate highly toxic continuations. We perform toxicity evaluation using the mean toxicity scores from the Perspective API [17] (appendix D).
@@
Table 1 shows results for the harmful QA task. HFAs without any defence mechanism substantially increase the harmfulness score (Base) on the base model. Attacks with higher learning rates and more data tend to be stronger. This replicates previous results about the effectiveness of HFAs at circumventing safety training in LLMs [8, 12, 29, 3, 13]. Evaluating our defence mechanisms, Security Vectors provides some resistance but RepNoise is the only defence method to consistently able to provide significant resistance across all attacks (Mann-Whitney $U$-test, $p<0.001$). [^2] We note that stronger attacks with thorough hyperparameter search can still succeed against RepNoise which we discuss in Appendix E.1).
@@ 4.3 Trainability
Model ViGGO E2E NLG DART CACAPO ConvWeather
@@
Table 4: ROUGE-1 score of RepNoise on GEM structured generation tasks before/after being fine-tuned.
@@ 4.4 Generalization
The results in Table 5 show that a defence using RepNoise is able to generalize to a defence against HFAs performed with unseen samples and unseen types of harm.

### v2

# Representation noising prevents harmful fine-tuning on LLMs

@@ Abstract
Releasing open-source large language models (LLMs) presents a dual-use risk since bad actors can easily fine-tune these models for harmful purposes. Even without the open release of weights, weight stealing and fine-tuning APIs make closed models vulnerable to harmful fine-tuning attacks (HFAs). While safety measures like preventing jailbreaks and improving safety guardrails are important, such measures can easily be reversed through fine-tuning. In this work, we propose Representation Noising (RepNoise), a defence mechanism that is effective even when attackers have access to the weights. RepNoise works by removing information about harmful representations such that it is difficult to recover them during fine-tuning. Importantly, our defence is also able to generalize across different subsets of harm that have not been seen during the defence process as long as they are drawn from the same distribution of the attack set. Our method does not degrade the general capability of LLMs and retains the ability to train the model on harmless tasks. We provide empirical evidence that the effectiveness of our defence lies in its “depth”: the degree to which information about harmful representations is removed across all layers of the LLM.
@@ 1 Introduction
Despite the benefits to both research and commercial development, open-sourcing large language models (LLMs) poses several risks [12] such as facilitating the development of weapons [25]. Such risks are not isolated to only open-source models, weights of proprietary models expose fine-tuning APIs [54] which can be used for constructing harmful models and for reconstructing weights at inference time [10]. The risk of LLMs assisting in harmful tasks is exacerbated by their increasing ability to follow instructions, carry out sophisticated tasks, and the ease with which they can be trained and run. Developers attempt to mitigate these risks [60] by developing safety guardrails that prevent LLMs from performing harmful tasks at inference time. However, these guardrails are easily circumvented either through back doors [37], adversarial attacks [50], or harmful fine-tuning [56]. We argue that no matter how sophisticated safety guardrails become, models vulnerable to harmful fine-tuning and amenable to malicious modifications are fundamentally unsafe.
We propose Representation Noising (RepNoise) as the first effective defence against in-distribution harmful fine-tuning attacks (HFAs) for LLMs in the natural language generation setting where the defender has no control of the model after the attacker attains its weights. Our work is inspired by the observation that safety mechanisms in LLMs are concentrated in a small proportion of the model weights (identified through ablation studies in [70]) and displace rather than replace harmful capabilities (identified by probing studies in [39, 42]); despite showing safe behaviour at inference-time, harmful behaviour can be easily recovered [56]. RepNoise works by removing the information structure of harmful representations such that they are much harder to recover during subsequent HFAs (fig. 1). We refer readers to § A.1 for precise definitions of representations, information in representations, and removing information.
@@
    We present extensive experimental evidence that our method prevents training on harmful question-answering and toxic content generation tasks while maintaining the ability to train the model on harmless tasks and preserving LLM capability (§ 4).
@@
Figure 1: Representation Noising pushes the intermediate activations of harmful text inputs (their representations) towards random directions, effectively reducing the mutual information between harmful representations and harmful text sequences and making it difficult to recover harmful representations through HFAs. We visualize this here as a projection (PCA) which isn’t able to recover any structure.
@@ 2 Harmful Fine-tuning and Defence Criteria
Harmful fine-tuning vulnerabilities in LLMs are established in several works [72, 44, 73, 56]. To formalize the problem, we borrow from Rosati et al.’s [58] harmful fine-tuning attack threat model.
@@
Harmful Fine-tuning Success: A HFA is formally designated a success if it causes the subject model to exceed a chosen threshold $\phi$ set by the defender on a given safety metric $g(\cdot):g(M_{\theta})>\phi$. Examples of safety metrics include Attack Success Rate [50, ASR] and probability outputs from the Jigsaw Perspective API [43]. In this formalism, a viable defence is one in which the attacker will have to spend more training steps $t$ to achieve $g(M_{\theta})>\phi$ than they can afford in time and/or compute.
@@
    Resistance: To increase the required effort for the attacker, a defended model $M_{\theta}^{*}$ should maximize the minimum number of training steps needed to produce a successful fine-tuning attack, i.e. $\max_{\theta}t(\theta)$ such that $g(M^{*}_{\theta[t]},D_{\text{harmful}})>\phi$
@@
    Stability: A defended model should preserve helpful (or harmless) behaviour of the undefended model. We model this using a reference dataset or task $D_{\text{ref}}$ where we want the defended model $M^{*}_{\theta}$ to perform approximately the same on some task measured by a task capability metric $f(\cdot)$, i.e. $f(M_{\theta[t=0]},D_{\text{ref}})\approx f(M^{*}_{\theta[t=0]},D_{\text{ref}})$. For example this could be ROUGE-1.
@@
    Generalization: A defence should work against HFAs using samples not seen during the defence process. Given disjoint subsets $D_{\text{harm}},D_{\text{harm}}^{\prime}\subset D_{\text{harmful}}$ the defence procedure producing $M^{*}_{\theta}$ based only on $D_{\text{harm}}$ should be resistant to HFAs using $D_{\text{harm}}^{\prime}$.
@@
    where $t_{1}$ and $t_{2}$ are the training steps for training the defended and undefended model equal within some small tolerance $\epsilon$ and $f(\cdot)$ is the same as in Stability.
A model that fulfils these conditions is said to be “immunized”. § 4 provides an operationalization of these conditions. Below, we provide the a defence that fulfils all these conditions.
@@ 3 Method
We propose Representation Noising, a method which fulfils the Immunization Criteria by reducing the mutual information between intermediate activations of harmful text sequences (their representations) and harmful text inputs before the attacker gains access and performs an HFA. Recall that after the attacker has access, the defender cannot intervene. There is evidence that current safety methods only suppress or route around preserved harmful representations [70, 39, 42]. This allows harmful behaviour to be easily recovered through HFAs. RepNoise instead aims to remove information about harmful tasks from intermediate activations over harmful text squences, to make it difficult for the model to relearn such information in future. The formal definition of information removal removal is based on mutual information between intermediate activations and generative outputs based on those activations and is specified in § A.1 which we encourage review before proceeding. RepNoise consists of a three-part loss: (i) reduce the predictive information in the weights for generating harmful outputs, (ii) retain capabilities by minimizing loss on harmless inputs, and (iii) push harmful representations towards random noise to remove harmful information. Fig. 1 presents a high-level intuition of our method.
@@
Our goal is to derive a loss function which will minimize the likelihood of recovering the mutual information $I(Z_{harmful};Y_{harmful})$ which is a quantity that measure how effective intermediate activations (or representations) $Z_{harmful}$ of harmful input sequences $X_{harmful}$ are at predicting the output token distribution $Y_{harmful}$. We are motivated by the observation in [1] that the number of training steps taken to fine-tune a model trained on one source task to another target task is minimized by $M_{\theta[t^{*}]}$) can be modelled as a transition probability over paths (or training trajectories) in a loss landscape. Formally in the language modeling case, a task here is a token output distribution over some dataset $D$. The source task is our initial pre-training distribution and the target task is the generative distribution of harmful text tokens in $D_{\text{harmful}}$. The loss landscape is the space ($\mathbf{R^{|}\theta|}$) of the value of a loss function $\mathcal{L}_{D}$ at every value of $\theta$. Paths or training trajectories in this landscape are the sequence of parameter configurations $\theta_{i}$ during a training procedure.
@@
Note that this is a simplified approximation of Achille et al. [1]; see appendix A for details. As the defender, we are only able to influence the initial set of weights $\theta_{t=0}$. Therefore our goal is to find a way to modify the model weights so that we minimize eq. 2. Where we have:
@@
A full proof of this is given in appendix A. Based on information bottleneck theory [64], we view multiple-layer neural networks as consisting of an encoder which maps the inputs $X$ to representations $Z$ and a decoder which maps the learned $Z$ to outputs $Y$. From Wang et al. [69] theorem 2, we can minimize the mutual information $I(Z;Y)$ directly by performing gradient ascent, which would decrease both the static distance and the reachability condition eq. 2 (see appendix A). However, we only want to minimize the transition probability that minimizes loss on a harmful dataset $D_{\text{harmful}}$. Therefore we only want to consider minimizing the mutual information $I(Z_{harmful};Y_{{harmful}})$ where harmful text sequences $X_{harmful}$ are passed through the model to construct intermediate activations $Z_{harmful}$ which are subsequently used to generate the output tokens $Y_{harmful}$.
$\ell_{\text{ascent}}$ $=\mathbb{E}_{(X_{harmful},Y_{harmful})\sim D_{\text{harmful}}}\mathcal{L}(M_{\theta}(X_{harmful}),Y_{harmful}).$ (3)
@@
$\mathcal{L}_{\text{Adversarial}}$ $=\ell_{\text{stability}}-\beta\cdot\ell_{\text{ascent}},$ (4)
@@
As we see later (Section 5), simply minimizing adversarial loss does not effectively remove the ability to predict harmful text sequences from the activations over harmful text sequences. This is because performing gradient ascent only minimizes $I(Y;Z)$—the mutual information between the output and decoder layers. However, according to the data processing inequality [71], $I(Y;X)\leq I(Z;X)$ despite having low mutual information $I(Y;Z)$, the mutual information between the inputs and the encoder layers $I(Z;X)$ can still be high. Consequently, it is possible for representations to retain the ability to generate harmful token sequences.
The above data processing inequality implies that minimizing $I(X;Z)$ minimizes $I(X;Y)$. We can do this by minimizing the KL divergence between the distribution of harmful representations given harmful input token sequences $\mathcal{P}_{\theta[t=0]}(Z\,|\,X)$ and Gaussian noise $\mathcal{N}(0,\textbf{{I}})$: if these two distributions were the same then $Z$ would have no information about $X$ [71]. This yields the noise loss $\ell_{\text{noise}}=KL(p(Z\,|\,X)\>||\>\mathcal{N}(0,\textbf{{I}}))$. Combining this with eq. 4 using another regularization term $\beta$, we get the loss function for Representation Noising (RepNoise) which satisfies Theorem 1.
$\mathcal{L}_{\text{RepNoise}}$ $=\mathcal{L}_{\text{Adversarial}}+\alpha\cdot\ell_{\text{noise}}=\ell_{\text{stability}}+\alpha\cdot\ell_{\text{noise}}-\beta\cdot\ell_{\text{ascent}}.$ (5)
@@ 4.1 Resistance
Here we simulate an HFA on llama-2-7b-chat and measure the harmfulness of the models and a series of controls before and after these attacks. Appendix K reports similar experiments on llama-2-13b-chat and the safety-trained Qwen (0.5B to 7B) series of models. We perform HFAs in two domains: harmful question-answering and toxic content generation [^1] We experimented with malicious code generation tasks [9] but observed base models did not guard against malicious code generation to begin with which is a pre-condition of performing our defence.. We measure attack strength in terms of the learning rate and number of samples used during supervised fine-tuning. Full details on our attack settings including rationale on learning rate choice can be found in appendix C.
To fine-tune for harmful question-answering, we use the BeaverTails harmful QA dataset [40] since it is a very large-scale dataset used in other attack literature [36], where the goal is to train an LLM to generate compliant answers to questions belonging to 14 categories of harm such as animal abuse and violent crime. For harmfulness evaluation, we use the logits of the harmful label after passing a question-answer pair into the harmfulness classifier trained on the BeaverTails dataset. The scores are computed as the mean of each individual logit score, for more details on how the classifier was trained as well as the scores is computed see § D.1.1. For toxic content generation, we use the DecodingTrust [67] split of Real Toxicity Prompts (RTP) [21] to fine-tune an LLM to generate highly toxic continuations. We perform toxicity evaluation using the mean toxicity scores from the Perspective API [43] (appendix D).
@@
Table 1 shows results for the harmful QA task. HFAs without any defence mechanism substantially increase the harmfulness score (Base) on the base model. Attacks with higher learning rates and more data tend to be stronger. This replicates previous results about the effectiveness of HFAs at circumventing safety training in LLMs [56, 72, 8, 54, 44]. Evaluating our defence mechanisms, Security Vectors provides some resistance but RepNoise is the only defence method to consistently able to provide significant resistance across all attacks (Mann-Whitney $U$-test, $p<0.001$). [^2] We note that stronger attacks with thorough hyperparameter search can still succeed against RepNoise which we discuss in Appendix E.1.
@@
Security Vectors 0.17 0.16 0.36 0.35
Vaccine ($\rho=1$) 0.19 0.46 0.70 0.72
Gradient Ascent 0.05 0.12 0.44 0.76
Adversarial loss 0.00 0.00 0.77 0.78
@@ 4.3 Trainability
Recall that Trainability is the defence condition from above that states that after applying defences models should still be able to be trained effectively on harmless datasets. The reason for this is that defences which remove or degrade training on harmless datasets are less useful than ones that do not under our threat model where defenders want to release these models such that they can still be trained on harmless tasks.
@@
ViGGO E2E NLG DART CACAPO ConvWeather
ROUGE-1
@@
Harmfulness
Base 0.03/0.75 0.05/0.65 0.05/0.69 0.06/0.67 0.05/0.55
RepNoise 0.00/0.00 0.16/0.01 0.00/0.00 0.02/0.27 0.01/0.08
Table 4: ROUGE-1 score of RepNoise on GEM structured generation tasks before/after being fine-tuned. Harmfulness scores before and after performing an attack at learning rate $3\times 10^{-5}$ with 1k samples from BeaverTails.
We further evaluated whether fine-tuning on a harmless task results in undoing safety guards or makes models more susceptible to HFAs. After fine-tuning on each GEM dataset, a HFA is performed with $3\times 10^{-5}$ with 1k samples from BeaverTails as above. Unlike the results of Qi et al. [56], both the base model and RepNoise are not made more harmful after harmless fine-tuning on GEM. However, training on GEM does seem to make the HFA more effective (readers can compare with the same attack in table 1). Even for RepNoise we see a small increase in attack efficacy after training the model on CACAPO which indicates the possibility that additional harmless fine-tuning could undo the RepNoise defence: a vulnerability which future work should explore.
We replicated Qi et al. [56] results in § E.3 (which RepNoise still provides an effective defence for) and the primary difference is that they are using a general instruction-following dataset rather than a specific task like structured generation to fine-tune the models.
@@ 4.4 Generalization
The results in Table 5 show that a defence using RepNoise is able to generalize to a defence against HFAs performed with unseen samples and unseen types of harm. However, it is important to note that these attacks are still in-distribution since the unseen types of harm are still drawn from the same BeaverTails distribution that the defender has seen. Importantly, RepNoise is not an effective defence against unseen sample out-of-distribution which we demonstrate in § E.2 using a distribution shift in the attack set to the HEX-PHI attack set [56].